Instrument Log

60M Chilled Mirror               C1 Lamont, OK

Date

Operator

Comments

8/7/01-8/8/01

MR

Southeast booms lowered at 1918 GMT on 8/7 for 25M and 60M T/H calibrations.  Booms were left down at ground level overnight.  Raised booms and RTS at 1840 GMT on 8/8.

8/7/01

MR

Tower systems is on site to perform maintenance on the 25M elevator and the 60M tower. Lowered the south east 25M elevator.  Tower systems replaced one of the connectors on the tower and performed general maintenance on the elevator and tower components.  RTS at 1910 GMT.

12/29/00

CAW

The 60 meter CHM needs removed from the tower and stored at the CF. Removed the CHM 1t 1445 GMT and gave to George Sawyer to store until it is needed again.

11/27/00

CAW

Lowered the SE 25m and 60m booms from 1840 to 2035 GMT to clean the chilled mirrors.

9/21/2000

CAM

David Cook requested that the SE 25 and 60 meter booms be lowered and reseated as part of troubleshooting bad humidity sensor readings. Lowered the booms slightly at 1850 GMT. Raised the booms and RTS at 1852 GMT.

5/1/98

CM/MW

Instrument mentor requested that the 60M Chilled Mirror be removed from service. Mentor needed the data logger returned to him for other use. Lowered 60M boom at 1826 GMT. Removed sensor, electronics box, wiring, and mounting hardware. Data logger will be shipped to mentor and the remaining components will be stored on site for future deployment. Raised boom at 1935 GMT.

12/19/97

CAM

Lowered booms from 2100 GMT to 2200 GMT to R/R SC32A with new SC32A supplied by mentor.

11/13/97

MW

SDS ops reported a loss of communications with the 60M chilled mirror instrument. The 60M GFCI was tripped. Reset the GFCI at 1728 GMT. RTS.